Sorry i don't know if this is under the right topic.
If a drink and sandwich together cost $5 and the sandwich costs 60c more than the drink how much does the drink cost.
i worked it out by
5-.60=$4.40
4.40/2=$2.20
so the drink costs $2.20 and the sandwich $2.80.
but how do i work this out using an algebric equation?
im sure its simple but i just can't get it.
thank you.

Answer by mananth(16949)

If a drink and sandwich together cost $5 and the sandwich costs 60c more than the drink how much does the drink cost.

Let the cost of drink be x cents
the sandwich costs 60 cents more than the drink
so the drink costs x+60 cents
x+x+60 = 500 cents
2x+60 = 500
2x=500-60
2x=440
x= 220 cents or $2.20 for the drink
So the sandwich costs $2.20 + 60 = $2.80